# listing[Listing][List of Listings]

### v1

@@ C.2 Standardized Prompts for Task 1
{listing}
[tb]Standardized Prompts for Task 1Role Setting: You are a large language model specialized in legal document analysis and information extraction, particularly skilled at identifying legal facts related to insider trading. Please read the following case description and extract the specified fields, outputting a standardized structured JSON.
Task Introduction: (1) Please strictly extract only the fields listed under the “Field Extraction Scope" below; (2) If a field is not mentioned in the Case Description, leave it as an empty string ‘ “" ’; (3) The output must be a valid JSON object, and the field names and nested structure must exactly follow the “Output Format Template”; (4) Do not generate any explanations or comments—output only the JSON data.
Case Description:
Field Extraction Scope:
Output Format Template: … {see Listing D “Structured Data Annotation Template" }
@@ C.3.1 Content of Insider Information
{listing}
[tb]Prompt for Content of Insider Information FieldPlease summarize the core facts of the insider information in 1–2 sentences using formal legal language. For example: “The major asset restructuring information of the company was leaked before public disclosure …”
@@ C.3.2 Type of Party Involved
{listing}
[tb]Prompt for Type of Party Involved FieldAccording to Article 51 of the Securities Law and Article 6 of the CSRC’s Guidelines for Identifying Insider Trading, determine the category of insider the involved party belongs to. Multiple selections are allowed from the following candidate list:
“51(1) Issuer and its directors, supervisors, senior executives”;
“Guideline 6(2-1) Issuer or listed company”;
“Guideline 6(2-2) Other companies controlled by controlling shareholders of the issuer”;
@@ C.3.3 Legal Provisions Applied
{listing}
[tb]Prompt for Legal Provisions Applied FieldPlease extract the legally effective article names and numbers from the following legal description. Only consider provisions from formal legal documents such as the Securities Law and Criminal Law. For example: “Article 73 of the Securities Law"; …
@@ C.3.4 Summary of Objections or Defenses
{listing}
[tb]Prompt for Summary of Objections or Defenses FieldOnly extract the objections or defenses claimed by the involved party, excluding comments from the court or regulatory authorities. Begin the summary with “The party argued that …”.
@@ Appendix E Field Categorization Details
Exa-Fields Insider Information Determination Name of the Traded Stock
Date of Insider Information Formation
@@
Basic Information of the Party Name
Gender
@@
Insider Trading Determination for the Party Date of Awareness
@@
Trade Amount (RMB)
@@
Sem-Fields Insider Information Determination Content of Insider Information
@@
Type of Insider Information
@@
Insider Trading Determination for the Party Role of the Party
@@
Manner in Which Insider Information Was Obtained
@@ Appendix G Case Study for Task 2
Figure 6: Reasoning results under Standard input.
Figure 7: Reasoning results under Human-Designed CoT input.

### v3

@@ C.2 Standardized Prompts for Task 1
Listing C.1: Standardized Prompts for Task 1 Role Setting: You are a large language model specialized in legal document analysis and information extraction, particularly skilled at identifying legal facts related to insider trading. Please read the following case description and extract the specified fields, outputting a standardized structured JSON. Task Introduction: (1) Please strictly extract only the fields listed under the “Field Extraction Scope” below; (2) If a field is not mentioned in the Case Description, leave it as an empty string “”; (3) The output must be a valid JSON object, and the field names and nested structure must exactly follow the “Output Format Template”; (4) Do not generate any explanations or comments—output only the JSON data. Case Description: Field Extraction Scope: Output Format Template: … (see Listing D “Structured Data Annotation Template”)
@@ C.3.1 Content of Insider Information
Listing C.2: Prompt for Content of Insider Information Field Please summarize the core facts of the insider information in 1–2 sentences using formal legal language. For example: “The major asset restructuring information of the company was leaked before public disclosure …”
@@ C.3.2 Type of Party Involved
Listing C.3: Prompt for Type of Party Involved Field According to Article 51 of the Securities Law and Article 6 of the CSRC’s Guidelines for Identifying Insider Trading, determine the category of insider the involved party belongs to. Multiple selections are allowed from the following candidate list: “51(1) Issuer and its directors, supervisors, senior executives”; “Guideline 6(2-1) Issuer or listed company”; “Guideline 6(2-2) Other companies controlled by controlling shareholders of the issuer”;
@@ C.3.3 Legal Provisions Applied
Listing C.4: Prompt for Legal Provisions Applied Field Please extract the legally effective article names and numbers from the following legal description. Only consider provisions from formal legal documents such as the Securities Law and Criminal Law. For example: “Article 73 of the Securities Law”; …
@@ C.3.4 Summary of Objections or Defenses
Listing C.5: Prompt for Summary of Objections or Defenses Field Only extract the objections or defenses claimed by the involved party, excluding comments from the court or regulatory authorities. Begin the summary with “The party argued that …”.
@@ Appendix E Field Categorization Details
Exa-Fields Name of the Traded Stock
Insider Information Determination Date of Insider Information Formation
@@
Name
Basic Information of the Party Gender
@@
Date of Awareness
@@
Insider Trading Determination for the Party Trade Amount (RMB)
@@
Sem-Fields Content of Insider Information
@@
Insider Information Determination Type of Insider Information
@@
Role of the Party
@@
Insider Trading Determination for the Party Manner in Which Insider Information Was Obtained
@@ Appendix G Case Study for Task 2
Figure 6: Reasoning results under Standard input.
Figure 7: Reasoning results under Human-Designed CoT input.